# Interpretable global minima of deep ReLU neural networks on sequentially separable data

### v1

@@ 1. Introduction
$\tau_{W,b}(x)=(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (1)
@@
$\left\{(\theta_{\ell},\nu_{\ell},\mu_{\ell})\right\}_{\ell=1}^{Q}\subset(0,\pi)\times\mathbb{R}^{M}\times(0,1),$ (3)
@@ 1.2. Outline of paper
In section 2, we clearly define the types of neural networks and the cost function under consideration, and fix notations. In section 3, we define the truncation map, remark on its action in terms of cones in input space and cumulative parameters, and state and prove key lemmas used in the proofs of the main results. In section 4, we give precise statements and proofs for proposition 4.1 and a corollary. In section 5, we define sequentially linearly separable data, and give a precise statement and proof for theorem 5.2. Finally, we collect some basic results on high dimensional geometry in appendix A.
@@ 2. Setting and notation
$x^{(\ell)}$ $=\sigma(W_{\ell}x^{(\ell-1)}+b_{\ell})\in\mathbb{R}^{d_{\ell}},\text{ for }\ell=1,\cdots,L-1,$ (4)
@@
$X^{(\ell)}=\sigma\left(W_{\ell}X^{(\ell-1)}+B_{\ell}\right),$ (5)
@@
$\overline{X_{0,j}}$ $:=X_{0,j}-\Delta X_{0,j}$ $\in\mathbb{R}^{M\times N_{j}},$ (10)
@@
$(\overline{X_{0}^{red}}){}^{+}=(\overline{X_{0}^{red}}^{T}\overline{X_{0}^{red}})^{-1}\overline{X_{0}^{red}}^{T},$ (11)
@@
$\left\lVert{A}\right\rVert_{\mathcal{L}^{2}}:=\sqrt{\operatorname{Tr}(AA^{T})},$ (13)
@@
$=\frac{1}{2}\sum_{j=1}^{Q}\sum_{i=1}^{N_{j}}\left\lvert{x_{0,j,i}^{(L)}-y_{j}}\right\rvert^{2}.$ (14)
@@
$=\left\lVert{X^{(L)}-Y^{ext}}\right\rVert_{\mathcal{L}_{\mathcal{N}}^{2}},$ (16)
@@
$=\sum_{j=1}^{Q}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{x_{0,j,i}^{(L)}-\overline{x_{0,j}^{(L)}}}\right\rvert^{2}+\left\lvert{\overline{x_{0,j}^{(L)}}-y_{j}}\right\rvert^{2}\right)$ (17)
$=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{\Delta x_{0,j,i}^{(L)}}\right\rvert^{2}+\sum_{j=1}^{Q}\left\lvert{\overline{x_{0,j}^{(L)}}-y_{j}}\right\rvert^{2},$
and since both terms are non-negative, the cost will be zero if, and only if, each term is zero. The global minimum with zero loss does not depend on the choice of weights for the cost function, and so we have the following interpretation for the results presented in this paper: The truncation maps implemented by the hidden layers have the role of minimizing the first term, by which all variances are reduced to zero, while the last (affine) layer minimizes the second term, whereby the class averages are matched with the reference outputs $y_{j}$.
@@ 3. Truncation map, cones and cumulative parameters
$x$ $\mapsto(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (18)
@@
$\tau_{W,b}(X)=(W){}^{+}\left(\sigma(WX+B)-B\right).$ (19)
@@
$W^{(\ell)}:=W_{\ell}\cdots W_{1},\quad\text{ for }\ell=1,\cdots,L,$ (21)
@@
$b^{(\ell)}:=\begin{cases}W_{\ell}\cdots W_{2}b_{1}+W_{\ell}\cdots W_{3}b_{2}+\cdots+W_{\ell}b_{\ell-1}+b_{\ell},&\text{ if }\ell\geq 2,\\ b_{1},&\text{ if }\ell=1,\end{cases}$ (22)
@@
$=W_{\ell}(W_{\ell}){}^{+}\left(\sigma(W_{\ell}X^{(\ell-1)}+B_{\ell})-B_{\ell}\right)+B_{\ell}$ (26)
@@
which proves (20). Going one step further,
@@
=$\stackrel{{\scriptstyle(*)}}{{=}}W_{\ell}W_{\ell-1}(W_{\ell}W_{\ell-1}){}^{+}\sigma(W_{\ell}(W_{\ell-1}\,\tau_{W_{\ell-1},b_{\ell-1}}(X^{(\ell-2)})+B_{\ell-1})+B_{\ell})$ (27)
@@
where ()(*)( ∗ ) follows from the surjectivity of $W_{\ell}W_{\ell-1}$. An iteration of these arguments then yields (23). Finally,
@@
This lemma, and especially expression (24), suggest that we think of a neural network with $L$ layers as a concatenation of $L-1$ truncation maps acting on the data in input space $\mathbb{R}^{d_{0}}$, followed by an affine map to the appropriate output space $\mathbb{R}^{d_{L}}$. These truncation maps and the affine map are given by cumulative parameters, so that from now on in this paper a layer will be defined by prescribing $(W^{(\ell)}$, $b^{(\ell)})$ instead of $(W_{\ell},b_{\ell})$, for all $\ell=1\cdots,L$.
@@
Consider a collection $\left\{W_{\ell}\right\}_{\ell=1}^{L}$ of surjective matrices, and assume the cumulative matrices $\{W^{(\ell)}\}_{\ell=1}^{L}$ defined as in (21) are also surjective. Define
@@
$\mathbb{R}_{+}^{n}:=\left\{(x_{1},\cdots,x_{n})^{T}\in\mathbb{R}^{n}:x_{i}\geq 0,i=1,\cdots,n\right\}$ (34)
@@
$\mathfrak{C}_{\theta}[h]:=\left\{x\in\mathbb{R}^{n}:\angle(x,h)\leq\frac{\theta}{2}\right\}.$ (35)
@@
$\theta_{n}:=2\arccos\frac{\sqrt{n-1}}{\sqrt{n}},$ (36)
@@
$\lambda(\theta,\theta_{n}):=\begin{cases}\frac{\tan\frac{\theta_{n}}{2}}{\tan\frac{\theta}{2}}<1,&\text{ if }\theta>\theta_{n},\\ 1,&\text{ otherwise.}\end{cases}$ (37)
@@
$\mathcal{I}(n,m)(e^{n}_{i})$ =$=\begin{cases}e^{m}_{i},\quad i=1,\cdots,m,\\ 0,\quad\text{else,}\end{cases}$ (38)
@@
$W_{\theta}=\tilde{R}\,\operatorname{diag}(1,\lambda(\theta,\theta_{n}),\cdots,\lambda(\theta,\theta_{n}))\,\tilde{R}^{T},$ (40)
with $\lambda(\theta,\theta_{n})$ defined as in (37) and $R,\tilde{R}\in O(n,\mathbb{R})$ are such that
$Rh=\frac{u_{n}}{\left\lvert{u_{n}}\right\rvert},\,\,\tilde{R}e^{n}_{1}=\frac{u_{n}}{\left\lvert{u_{n}}\right\rvert}.$ (41)
@@
$\tau_{W,b}(x)=\begin{cases}((W){}^{+}W)\,x,\quad x\in p+\mathfrak{C}_{\theta}[h],\\ ((W){}^{+}W)\,p,\quad x\in p+\mathfrak{C}_{\theta}[-h].\end{cases}$ (42)
@@
$\tau_{W,b}(x)=(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (44)
@@
$R(x-p)\in\mathfrak{C}_{\theta}[Rh]=\mathfrak{C}_{\theta}[u_{n}].$ (46)
@@
$W_{\theta}R(\mathfrak{C}_{\theta}[h])\subset\mathfrak{C}_{\theta_{n}}[u_{n}]\subset\mathbb{R}_{+}^{n},$ (47)
@@
$W(\mathfrak{C}_{\theta}[-h])\subset\mathfrak{C}_{\theta_{n}}[-u_{n}]\subset\mathbb{R}_{-}^{n},$ (48)
@@
$\tau_{W,b}(x)=(W){}^{+}(-b)=(W){}^{+}Wp.$ (49)
Finally, (43) follows easily from the definition of the truncation map,
@@
$=(W){}^{+}\left(\sigma(Wx+b)-b\right)$ (50)
@@
$W=\left\lvert{W}\right\rvert R$ (52)
@@
$=\tilde{R}\operatorname{diag}(1,\cdots)^{2}\tilde{R}^{T}$ (53)
@@ 4. Clustered data and dimensional reduction
$\delta:=\sup_{i,j}\left\lvert{\Delta x_{0,j,i}}\right\rvert,$ (54)
@@
$\delta<c_{0}\min_{j}\left\lvert{\overline{x_{0,j}}-\overline{x}}\right\rvert,$ (55)
@@
$\overline{x}:=\frac{1}{Q}\sum_{j=1}^{Q}\overline{x_{0,j}},$ (56)
@@
$f_{j}:=\frac{\overline{x}-\overline{x_{0,j}}}{\left\lvert{\overline{x}-\overline{x_{0,j}}}\right\rvert}$ (57)
@@
$\bigcup_{j^{\prime}\neq j}B_{4\delta}(\overline{x_{0,j^{\prime}}})\subset\overline{x_{0,j}}+\mathfrak{C}_{\theta_{*,j}}[f_{j}],$ (58)
@@
Consider a set of training data $\mathcal{X}_{0}=\bigcup_{j=1}^{Q}\mathcal{X}_{0,j}\subset\mathbb{R}^{M}$ separated into $Q$ classes corresponding to linearly independent labels $\{y_{j}\}_{j=1}^{Q}\subset\mathbb{R}^{Q}$. If conditions (55) and (59) are satisfied, then a neural network with ReLU activation function defined as in (2), with $L=Q+1$ layers, $d_{0}=M$, $d_{Q+1}=Q$, and $d_{\ell-1}\geq d_{\ell}\geq Q$ for all $\ell=1,\cdots,L$, attains
$\min_{(W_{i},b_{i})_{i=1}^{L}}\mathcal{C}[(W_{i},b_{i})_{i=1}^{L}]=0,$ (60)
with weights and biases given recursively by (21), (22) and the following cumulative parameters: for $\ell=1,\cdots,Q$,
@@
$x=\sum_{j=1}^{Q}\kappa_{j}\overline{x_{0,j}}+\tilde{x}$ (63)

### v2

@@ 1. Introduction
$\tau_{W,b}(x)=(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (1.1)
@@
$\left\{(\theta_{\ell},\nu_{\ell},\mu_{\ell})\right\}_{\ell=1}^{Q}\subset(0,\pi)\times\mathbb{R}^{M}\times(0,1),$ (1.3)
@@ 1.1. Related work
Finally, the results here show how a trained neural network could take within-class variability to zero. This is one of the features of neural collapse, a phenomenon which has been empirically shown in [PHD20, HPD22] to emerge when training a neural network towards zero loss.
@@ 1.2. Outline of paper
In section 2, we clearly define the types of neural networks and the cost function under consideration and fix notation. In section 3, we define the truncation map, remark on its action in terms of cones in input space and cumulative parameters, and prove key lemmas. In section 4, we give precise statements for Proposition 4.1 and a corollary. In section 5, we define sequentially linearly separable data, and give a precise statement and proof for Theorem 5.2. In section 6, we prove Proposition 4.1. Finally, we collect some basic results on geometry of cones in appendix A.
@@ 2. Setting and notation
$x^{(\ell)}$ $=\sigma(W_{\ell}x^{(\ell-1)}+b_{\ell})\in\mathbb{R}^{d_{\ell}},\text{ for }\ell=1,\cdots,L-1,$ (2.1)
@@
$X^{(\ell)}=\sigma\left(W_{\ell}X^{(\ell-1)}+B_{\ell}\right),$ (2.2)
@@
$\overline{X_{0,j}}$ $:=X_{0,j}-\Delta X_{0,j}$ $\in\mathbb{R}^{M\times N_{j}},$ (2.7)
@@
$(\overline{X_{0}^{red}}){}^{+}=(\overline{X_{0}^{red}}^{T}\overline{X_{0}^{red}})^{-1}\overline{X_{0}^{red}}^{T},$ (2.8)
@@
$\left\lVert{A}\right\rVert_{\mathcal{L}^{2}}:=\sqrt{\operatorname{Tr}(AA^{T})},$ (2.10)
@@
$=\frac{1}{2}\sum_{j=1}^{Q}\sum_{i=1}^{N_{j}}\left\lvert{x_{0,j,i}^{(L)}-y_{j}}\right\rvert^{2}.$ (2.11)
@@
$=\left\lVert{X^{(L)}-Y^{ext}}\right\rVert_{\mathcal{L}_{\mathcal{N}}^{2}},$ (2.13)
@@
$=\sum_{j=1}^{Q}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{x_{0,j,i}^{(L)}-\overline{x_{0,j}^{(L)}}}\right\rvert^{2}+\left\lvert{\overline{x_{0,j}^{(L)}}-y_{j}}\right\rvert^{2}\right)$ (2.14)
$=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{\Delta x_{0,j,i}^{(L)}}\right\rvert^{2}+\sum_{j=1}^{Q}\left\lvert{\overline{x_{0,j}^{(L)}}-y_{j}}\right\rvert^{2}$
$=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{W^{(L)}\Delta x_{0,j,i}^{(\tau,L-1)}}\right\rvert^{2}+\sum_{j=1}^{Q}\left\lvert{W^{(L)}\overline{x_{0,j}^{(\tau,L-1)}}+b^{(L)}-y_{j}}\right\rvert^{2},$
and since both terms are non-negative, the cost will be zero if, and only if, each term is zero. The global minimum with zero loss does not depend on the choice of weights for the cost function, [^4] That is, the constants in $\mathcal{N}$ could be any sequence $(\alpha_{j})_{j=1}^{Q}$; common choices are $\alpha_{j}=1$ or $\alpha_{j}=\frac{1}{2}$ for all $j=1,\cdots,Q$. and so we have the following interpretation for the results presented in this paper: The truncation maps implemented by the hidden layers have the role of minimizing the first term, by which all variances are reduced to zero, while the last (affine) layer minimizes the second term, whereby the class averages are matched with the reference outputs $y_{j}$.
@@ 3. Truncation map, cones and cumulative parameters
$x$ $\mapsto(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (3.1)
@@
$\tau_{W,b}(X)=(W){}^{+}\left(\sigma(WX+B)-B\right).$ (3.2)
@@
$W^{(\ell)}:=W_{\ell}\cdots W_{1},\quad\text{ for }\ell=1,\cdots,L,$ (3.4)
@@
$b^{(\ell)}:=\begin{cases}W_{\ell}\cdots W_{2}b_{1}+W_{\ell}\cdots W_{3}b_{2}+\cdots+W_{\ell}b_{\ell-1}+b_{\ell},&\text{ if }\ell\geq 2,\\ b_{1},&\text{ if }\ell=1,\end{cases}$ (3.5)
@@
$=W_{\ell}(W_{\ell}){}^{+}\left(\sigma(W_{\ell}X^{(\ell-1)}+B_{\ell})-B_{\ell}\right)+B_{\ell}$ (3.9)
@@
which proves (3.3). Going one step further,
@@
=$\stackrel{{\scriptstyle(*)}}{{=}}W_{\ell}W_{\ell-1}(W_{\ell}W_{\ell-1}){}^{+}\sigma(W_{\ell}(W_{\ell-1}\,\tau_{W_{\ell-1},b_{\ell-1}}(X^{(\ell-2)})+B_{\ell-1})+B_{\ell})$ (3.10)
@@
where ()(*)( ∗ ) follows from the surjectivity of $W_{\ell}W_{\ell-1}$. An iteration of these arguments then yields (3.6). Finally,
@@
This lemma, and especially expression (3.7), suggest that we think of a neural network with $L$ layers as a concatenation of $L-1$ truncation maps acting on the data in input space $\mathbb{R}^{d_{0}}$, followed by an affine map to the appropriate output space $\mathbb{R}^{d_{L}}$. These truncation maps and the affine map are given by cumulative parameters, so that from now on in this paper a layer will be defined by prescribing $(W^{(\ell)}$, $b^{(\ell)})$ instead of $(W_{\ell},b_{\ell})$, for all $\ell=1\cdots,L$.
@@
Consider a collection $\left\{W_{\ell}\right\}_{\ell=1}^{L}$ of surjective matrices, and assume the cumulative matrices $\{W^{(\ell)}\}_{\ell=1}^{L}$ defined as in (3.4) are also surjective. Define
@@
$\mathbb{R}_{+}^{n}:=\left\{(x_{1},\cdots,x_{n})^{T}\in\mathbb{R}^{n}:x_{i}\geq 0,i=1,\cdots,n\right\}$ (3.17)
@@
$\mathfrak{C}_{\theta}[h]:=\left\{x\in\mathbb{R}^{n}:\angle(x,h)\leq\frac{\theta}{2}\right\}.$ (3.18)
@@
$\theta_{n}:=2\arccos\frac{\sqrt{n-1}}{\sqrt{n}},$ (3.19)
@@
$\lambda(\theta,\theta_{n}):=\begin{cases}\frac{\tan\frac{\theta_{n}}{2}}{\tan\frac{\theta}{2}}<1,&\text{ if }\theta>\theta_{n},\\ 1,&\text{ otherwise.}\end{cases}$ (3.20)
@@
$\mathcal{I}(n,m)(e^{n}_{i})$ =$=\begin{cases}e^{m}_{i},\quad i=1,\cdots,m,\\ 0,\quad\text{else,}\end{cases}$ (3.21)
@@
$W_{\theta}=\tilde{R}\,\operatorname{diag}(1,\lambda(\theta,\theta_{n}),\cdots,\lambda(\theta,\theta_{n}))\,\tilde{R}^{T},$ (3.23)
with $\lambda(\theta,\theta_{n})$ defined as in (3.20) and $R,\tilde{R}\in O(n,\mathbb{R})$ are such that
$Rh=\frac{u_{n}}{\left\lvert{u_{n}}\right\rvert},\,\,\tilde{R}e^{n}_{1}=\frac{u_{n}}{\left\lvert{u_{n}}\right\rvert}.$ (3.24)
@@
$\tau_{W,b}(x)=\begin{cases}((W){}^{+}W)\,x,\quad x\in p+\mathfrak{C}_{\theta}[h],\\ ((W){}^{+}W)\,p,\quad x\in p+\mathfrak{C}_{\theta}[-h].\end{cases}$ (3.25)
@@
$\tau_{W,b}(x)=(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (3.27)
@@
$R(x-p)\in\mathfrak{C}_{\theta}[Rh]=\mathfrak{C}_{\theta}[u_{n}].$ (3.29)
@@
$W_{\theta}R(\mathfrak{C}_{\theta}[h])\subset\mathfrak{C}_{\theta_{n}}[u_{n}]\subset\mathbb{R}_{+}^{n},$ (3.30)
@@
$W(\mathfrak{C}_{\theta}[-h])\subset\mathfrak{C}_{\theta_{n}}[-u_{n}]\subset\mathbb{R}_{-}^{n},$ (3.31)
@@
$\tau_{W,b}(x)=(W){}^{+}(-b)=(W){}^{+}Wp.$ (3.32)
Finally, (3.26) follows easily from the definition of the truncation map,
@@
$=(W){}^{+}\left(\sigma(Wx+b)-b\right)$ (3.33)
@@
$W=\left\lvert{W}\right\rvert R$ (3.35)
@@
$=\tilde{R}\operatorname{diag}(1,\cdots)^{2}\tilde{R}^{T}$ (3.36)
@@ 4. Clustered data and dimensional reduction
$\delta:=\sup_{i,j}\left\lvert{\Delta x_{0,j,i}}\right\rvert,$ (4.1)
@@
$\delta<c_{0}\min_{j}\left\lvert{\overline{x_{0,j}}-\overline{x}}\right\rvert,$ (4.2)
@@
$\overline{x}:=\frac{1}{Q}\sum_{j=1}^{Q}\overline{x_{0,j}},$ (4.3)
@@
$f_{j}:=\frac{\overline{x}-\overline{x_{0,j}}}{\left\lvert{\overline{x}-\overline{x_{0,j}}}\right\rvert}$ (4.4)
@@
$\bigcup_{j^{\prime}\neq j}B_{4\delta}(\overline{x_{0,j^{\prime}}})\subset\overline{x_{0,j}}+\mathfrak{C}_{\theta_{*,j}}[f_{j}],$ (4.5)
@@
Consider a set of training data $\mathcal{X}_{0}=\bigcup_{j=1}^{Q}\mathcal{X}_{0,j}\subset\mathbb{R}^{M}$ separated into $Q$ classes corresponding to linearly independent labels $\{y_{j}\}_{j=1}^{Q}\subset\mathbb{R}^{Q}$. If conditions (4.2) and (4.6) are satisfied, then a neural network with ReLU activation function defined as in (2), with $L=Q+1$ layers, $d_{0}=M$, $d_{Q+1}=Q$, and $d_{\ell-1}\geq d_{\ell}\geq Q$ for all $\ell=1,\cdots,L$, attains
$\min_{(W_{i},b_{i})_{i=1}^{L}}\mathcal{C}[(W_{i},b_{i})_{i=1}^{L}]=0,$ (4.7)
with weights and biases given recursively by (3.4), (3.5) and the following cumulative parameters: for $\ell=1,\cdots,Q$,
@@
$x=\sum_{j=1}^{Q}\kappa_{j}\overline{x_{0,j}}+\tilde{x}$ (4.10)